rose is 25 years younger than her father. 5 years ago her father was twice as old as she will be 5 years from now. how old is rose today?
Answer by stanbon(75887)

Rose is 25 years younger than her father. 5 years ago her father was twice as old as she will be 5 years from now. how old is rose today?
--------------------
r = f - 25
f-5 = 2(r+5)
------------------------
Substitute for "r" and solve for "f"::
f-5 = 2(f-25)+10
------
f-5 = 2f - 40
---------------
f = 35 (father's age now)
r = f-25 = 10 (rose's age now)
---------------
Cheers,
Stan H.
----------------